You need a particular piece of equipment for your production process. An equipment-leasing company has offered to lease the equipment to you for $10,100 per year if you sign a guaranteed 5 year lease? (the lease is paid at the end of each? year). The company would also maintain the equipment for you as part of the lease.?Alternatively, you could buy and maintain the equipment yourself. The cash flows from doing so are listed below? (the equipment has an economic life of 5 years). If your discount rate is 7.4%, what should you? do?

Year 0 = $-40,200 Year 1-5 = $-1,800
Answer #1

Present value of leasing option

Present value is calculated as below :

Present value of each cash flow = cash flow / (1 + discount rate)number of years

Present value of leasing option is -$40,972.

C8 fi =SUM(C3:07) А 1 Year в Cash Flow PV vwN 0 1 2 3 4 5 ($10,100) ($10,100) ($10,100) ($10,100) ($10,100) ($9,404) ($8,756)

Cash Flow PV А 1 Year 2 0 31 42 5 3 64 7 5 -10100 -10100 -10100 -10100 -10100 =B3/(1+7.4%)^A3 =B4/(1+7.4%)^A4 =B5/(1+7.4%)^A5

Present value of buying option

Present value is calculated as below :

Present value of each cash flow = cash flow / (1 + discount rate)number of years

Present value of buying option is -$47,502.

019 - fx =SUM(C13:C18) A B C DE 12 Year Cash Flow PV 13 0 ($40,200) ($40,200) 1 ($1,800) ($1,676) 2 ($1,800) $1,561) 3 ($1,80

C PV А 12 Year 13 O 14 1 15 2 16 3 17 4 18 5 19 20 | B Cash Flow -40200 -1800 -1800 -1800 -1800 -1800 T=B13/(1+7.4%)^A13 =B14

The present value of leasing option is lower. Therefore you should lease.